OpenStudy (anonymous):

The probability of selecting a green marble at random from a jar that contains only green, white, and yellow marbles 1/4. The probability of selecting a white marble at random from the same jar is 1/3. If this jar contains 10 yellow marbles, what is the total number of marbles in the jar?

OpenStudy (kewlgeek555):

I believe it is twenty four.

OpenStudy (kewlgeek555):

No. 24 is the answer.

OpenStudy (kewlgeek555):

Okay...

OpenStudy (kewlgeek555):

The rest of the question?

OpenStudy (kewlgeek555):

Isn't it one question all in one, or is it two questions I am supposed to answer?

hartnn (hartnn):

the sum of probabilities add up to 1 so can u find the probability of selecting yellow marble??

OpenStudy (kewlgeek555):

Oh, I have choices? What are my choices?

OpenStudy (kewlgeek555):

okay.

OpenStudy (kewlgeek555):

I keep getting 24. Hmm.... I'll have to check the internet.

hartnn (hartnn):

@sakigirl can u find the probability of selecting yellow marble?

OpenStudy (kewlgeek555):

The internet says 24, too.

hartnn (hartnn):

ok,let me do it. p(green) + p(white) + p(yellow) =1 so p(yellow) = 1-1/3-1/4=5/12 now probability is defined by favorable events / total events so p(yellow) = no. of yellow balls / total balls total balls = no. of yellow balls/p(yellow) =10/(5/12) =10*12/5 =24